OpenStudy (anonymous):

**WILL MEDAL** Rudy delivers groceries from his uncle’s store to earn extra money. His uncle pays him $25.00 each day he works and an additional $2.50 for each delivery he makes. Which table shows values for the amount Rudy makes in a day based upon the number of deliveries he makes?

OpenStudy (anonymous):

A. Rudy’s Earnings Number of deliveries 2 4 7 9 Amount earned $30.00 $32.50 $42.50 $47.50 B. Rudy’s Earnings Number of deliveries 2 5 7 10 Amount earned $30.00 $37.50 $45.00 $50.00 C. Rudy’s Earnings Number of deliveries 3 5 8 10 Amount earned $32.50 $37.50 $45.00 $50.00 D. Rudy’s Earnings Number of deliveries 3 6 8 11 Amount earned $35.00 $42.50 $47.50 $55.00

OpenStudy (anx):

He's given $25 a day for his work.. and $2.50 for every delivery so your equation would look something like 25 + 2.50x = y Where x is the delievers and y is the total amount of money he makes. Just plugin the values your given in you "delivery column for x" and see if it matches the total amount (y) If you need further help, lemme know

OpenStudy (anonymous):

So B or C

OpenStudy (anx):

Let's evaluate (B) x = 2,5,7,10 25 + 2.50(2) = 30 YES 25+2.50(5) = 37.50 No 25+2.50(7) = 42.5 No 25+2.50(10) = 50 NO (C) x = 3,5,8,10 25 + 2.50(3) = $32.5 YES 25+2.50(5) = $37.50 YES 25+2.50(8) = $45 YES 25+2.50(10) = $50 YES so C would be your answer.

OpenStudy (anonymous):

Thx